OpenStudy (anonymous):

Select the inequality that corresponds to the given graph. number line with a closed circle on negative 2 and 5 with shading in between |2x - 4| > 2 |4x + 2| > 10 |4x - 6| less than or greater to 14 |6x + 4| less than or greater to 16

OpenStudy (anonymous):

@Derrick14 @surjithayer @sam801 @Luigi0210 @countonme123 @flyguy6 @Ryaan @genius12

OpenStudy (anonymous):

|4x - 6| <= 14 ** |6x + 4| <= 16 ***

OpenStudy (anonymous):

@Felicia123 @Hero @terenzreignz @Hero @keysf102

OpenStudy (anonymous):

there will not be any option unlees it is given less than or equal to

OpenStudy (anonymous):

|2x - 4| > 2 |4x + 2| > 10 |4x - 6| <= 14 |6x + 4| <=16 these are the options

OpenStudy (anonymous):

when the shaded region is in between the two always consider < or= so check only last two

OpenStudy (anonymous):

do you know what the answer Is?

OpenStudy (anonymous):

\[\left| 4x-6 \right|\le 14,-14\le4x-6\le 14\]

OpenStudy (anonymous):

add 6

OpenStudy (anonymous):

\[-14+6 \le 4x-6+6 \le 14+6\] \[-8 \le 4x \le 20\] divide by 4 \[-2 \le x \le 5\]

OpenStudy (anonymous):

what is the answer ? out of the choice that are up there

OpenStudy (anonymous):

you can clearly see it is third .

OpenStudy (anonymous):

are u sure?

OpenStudy (anonymous):

yes